# Principal stratification with recurrent events truncated by a terminal event: A nested Bayesian nonparametric approach

### v2

@@ 4 Simulation Studies
Table 2: Bias and root mean squared error (RMSE) of point estimates, and coverage probability (CP) and average length (AL) of $95\%$ central credible intervals of the key estimands.
$\mu^{0}_{500}(500)$ $\mu^{1}_{500}(500)$ $\tau^{0}_{500}(500)$ $\tau^{1}_{500}(500)$ Model Bias RMSE CP AL Bias RMSE CP AL Bias RMSE CP AL Bias RMSE CP AL LM 0.08 0.09 0.40 0.10 0.19 0.21 0.42 0.15 11.84 12.75 0.42 7.83 12.31 14.17 0.41 7.51 DPM 0.13 0.14 0.08 0.16 0.18 0.19 0.17 0.20 11.02 11.38 0.04 13.72 9.69 10.01 0.18 13.09 DDPM 0.09 0.09 0.29 0.14 0.19 0.20 0.06 0.19 10.41 10.84 0.09 13.50 9.79 10.10 0.07 11.83 EDDPM 0.02 0.03 0.97 0.15 0.06 0.08 0.80 0.20 -1.28 3.63 1.00 14.55 0.52 3.06 0.96 12.57
5 An Illustrative Data Example
We demonstrate our proposed methods by analyzing the SPRINT dataset introduced in Section 1.1. Our primary goals include evaluating the causal esimands on recurrent hospital visits defined in Section 2.2, i.e., differences between the treatment and control groups in the expected number of hospital visits (SANR) and the average event rate (SAER) for always-survivors. Specifically, we consider $\text{SANR}(t;r)=\mu^{1}\quantity(t;r)/\mu^{0}\quantity(t;r)$ and $\text{SAER}(t;r)=\quantity{\mu^{1}\quantity(t;r)/\tau^{1}\quantity(t;r)}/\quantity{\mu^{0}\quantity(t;r)/\tau^{0}\quantity(t;r)}$, for $t<r$. Understanding these aspects can clarify whether and how structured exercise modifies not only the frequency but also the timing patterns of hospital visits, which are critical dimensions of clinical management and quality of life in patients with chronic heart failure. We perform inference using model (7) with proper, weakly informative conjugate priors for all parameters. Specifically, the chosen priors are identical to those used in the simulation study presented in Section 4. We set $\rho=0.5$ as the default correlation parameter and examine the sensitivity for different values.
We evaluate the predictive performance of the models using the log pseudo marginal likelihood (LPML; Geisser1979). For the Bayesian models considered in simulations—LM, DPM, DDPM, and EDDPM—the LPML values are 21113.86-21113.86, 16397.04-16397.04, 15203.78-15203.78, and 12245.87-12245.87, respectively. Since a higher LPML indicates a better predictive fit in terms of LOO predictive densities, the EDDPM emerges as the superior model among the four models. We therefore focus on the EDDPM method in the subsequent discussion.
Figure 2: Contour plots of posterior mean (left), $95\%$ posterior lower bound (middle), and upper bound (right) for SANR. The posterior values are computed for different cut-off values of $t$ and $r$ from $400$ days ($\approx 1$ year) to $2000$ days ($\approx 5$ years) with an increment of $200$ such that $t\leq r$, and interpolated between the grid values with a cubic spline. The regions with the estimate greater than 11 are represented using warm colors (yellow to red), while regions with the estimate less than 11 are indicated using cool colors (blue).
Figure 3: Contour plots of posterior mean (left), $95\%$ posterior lower bound (middle), and upper bound (right) for SAER.
Figures 2 and 3 display contour plots of the posterior mean along with the corresponding $95\%$ credible lower and upper surfaces, where the horizontal axis tracks follow-up time $t$ and the vertical axis fixes the always-survivor subpopulation indexed by $r$. For SANR in Figure 2, the $95\%$ upper surface remains almost entirely below $1.0$ for $t>700$, providing evidence that an intensive SBP reduces the expected number of recurrent major cardiovascular events over the long term; the posterior distributions also declines with increasing $t$, suggesting that the treatment effect becomes more pronounced as follow-up progresses. Similarly, the $95\%$ upper surface of SAER in Figure 3 remains almost entirely below $1.0$ for $t>700$, showing evidence that the intensive SBP reduces the expected rate of recurrent major cardiovascular events over the long term. Across both estimands, treatment effects appear largely homogeneous with respect to $r$, indicating minimal effect-modification by this dimension.
To gauge the reliability of the Bayesian analyses for weakly identifiable models, we undertake a sensitivity analysis focusing on the unidentifiable frailty correlation parameter $\rho$. Specifically, we investigate how the posterior distribution of the SANR estimand responds to more extreme values of $\rho\in\{0.1,0.9\}$, which represent weak and strong correlation between cross-world frailties, and hence potential outcomes. Figure 4 displays the resulting posterior estimates. Switching from $\rho=0.5$ to the lower correlation $\rho=0.1$ increases the upper bounds by roughly 0.010.10.01-0.1 across all time points, whereas switching to $\rho=0.9$ shift the posterior distributions only slightly. Importantly, even under these slight shifts, the upper bounds of the credible intervals decline with increasing $t$ and remain below $1.0$ for $t>1000$ for both extreme scenarios, indicating that the substantive conclusions are unaffected, supporting the claim that intensive SBP control exerts a durable, clinically meaningful reduction in the expected number and rate of recurrent major cardiovascular events. Additional sensitivity analyses for additional values of $\rho$ and for the SAER are provided in Web Appendix E.
Figure 4: Posterior distributions of SANR under two alternative sensitivity parameters $\rho$, displayed vertically: top, $\rho=0.1$; bottom, $\rho=0.9$.
@@ 6 Concluding Remarks
We proposed a new principal stratification approach for causal recurrent event analysis in the presence of semi-competing risks. Our first significant contribution is the introduction of new principal causal estimands. We critically examine the limitations inherent in the traditional definition of always-survivors within recurrent event analyses and propose a refined definition that more accurately reflects treatment effects on recurrent events over time. Our second methodological innovation is the adaptation of the EDDPM prior, which incorporates a nested structure tailored for joint analyses of recurrent events with semi-competing risks and embeds covariate dependence into the prior. The EDDPM effectively captures distributional heterogeneity and the complex nested dependencies between recurrent and terminal events. Furthermore, it addresses and mitigates the unfavorable random partition structures induced by traditional DP priors, an issue commonly neglected in similar analyses. We provided a fully tractable Gibbs sampling algorithm for inference, along with a rigorous theoretical investigation into the random partition structure imposed by our proposed prior. Simulation studies demonstrate that our method significantly outperforms traditional DP-based methods in accuracy and robustness. The practical utility of our approach was further illustrated through an analysis of data from the SPRINT trial, a real-world randomized clinical trial.
We adopted a frailty approach to explicitly model the dependence between recurrent and terminal event processes. The frailty is modeled using our proposed nested EDP prior, which offers flexibility over restrictive parametric approaches (Nevo2022; Lyu2023) and resolves the undesirable random partitioning effects of existing works (e.g., Paulon2020). Alternatively, cross-world and within-world dependencies could also be modeled using copula-based methods, such as those proposed by Xu2022. Importantly, even within the copula framework, our nested prior remains potentially beneficial and applicable, offering advantages in terms of avoiding undesirable partition issues.

### v3

@@ 4 Simulation Studies
Table 4: Bias and root mean squared error (RMSE) of point estimates, and coverage probability (CP) and average length (AL) of $95\%$ central credible intervals of $\mu^{z}(t;r)$ for $(t,r)=(300,500),(500,500)$.
$\mu^{0}(t;r)$ $\mu^{1}(t;r)$ $(t,r)$ Model Bias RMSE CP AL Bias RMSE CP AL $(300,500)$ LM 0.05 0.09 0.23 0.12 0.04 0.08 0.18 0.09 DPM 0.08 0.08 0.78 0.17 0.01 0.06 0.45 0.09 DDPM 0.03 0.06 0.83 0.12 0.00 0.04 0.74 0.09 EDDPM 0.01 0.03 0.85 0.09 0.02 0.03 0.80 0.08 $(500,500)$ LM 0.09 0.15 0.16 0.18 0.06 0.12 0.10 0.13 DPM 0.11 0.12 0.92 0.30 -0.00 0.09 0.44 0.14 DDPM 0.03 0.08 0.84 0.18 -0.01 0.06 0.67 0.13 EDDPM -0.01 0.05 0.87 0.14 0.01 0.04 0.87 0.12
We compare our proposed EDDPM (5) with the models commonly used in the recurrent event analysis: linear model (LM) (e.g., Comment et al., 2025; Lyu et al., 2023), Dirichlet process mixture (DPM) (e.g., Paulon et al., 2020; Tian et al., 2024), and dependent Dirichlet process mixture (DDPM) (e.g., Xu et al., 2022). In DDPM, the nested prior structure is removed from EDDPM. In DPM, covariate dependence is further removed from DDPM, and the frailty and error terms are modeled using DPM, resulting in a model equivalent to that of Paulon et al. (2020). For LM, we fit a linear mixed model to the log-transformed survival and gap times. The specification includes a subject-level frailty term and assumes normal residual errors, which yields the familiar accelerated failure time (AFT) representation. This setup is close to the models explored by Comment et al. (2025) and Lyu et al. (2023) and is strictly parametric. The only difference is that their proportional hazard model assumes an extreme-value distribution for the error term. We use proper, weakly informative conjugate priors for all parameters. Specifically, we choose the hyperparameters such that $a_{\alpha_{\phi}}=a_{\alpha_{\theta\mid\phi}}=a_{\tau}=a_{\sigma}=2.0$, $b_{\alpha_{\phi}}=b_{\alpha_{\theta\mid\phi}}=b_{\tau}=b_{\sigma}=1.0$, $\mu_{\gamma}=\mu_{\psi}=0.0$, $\sigma^{2}_{\gamma}=\sigma^{2}_{\psi}=3.0^{2}$, $\boldsymbol{\mu}_{\boldsymbol{\beta}_{u}}$ and $\boldsymbol{\mu}_{\boldsymbol{\beta}_{y}}$ are zero-vectors of appropriate sizes, and $\boldsymbol{\Sigma}_{\boldsymbol{\beta}_{u}}$ and $\boldsymbol{\Sigma}_{\boldsymbol{\beta}_{y}}$ are diagonal matrices with $3.0^{2}$ on the diagonal elements.
Table 4 presents the simulation results. Overall, the results consistently demonstrate that our method yields unbiased estimates and has the smallest RMSE across all scenarios, indicating EDDPM’s superior accuracy in point estimation. Turning to interval estimation, which is assessed by the coverage probability (CP) and average length (AL) of $95\%$ credible intervals, EDDPM stands out for better CP compared to other methods, closer to the frequentist target $95\%$, and shorter interval lengths (reflecting improved efficiency). The linear model is sensitive to latent treatment heterogeneity within the data-generating processes, resulting in the poorest performance across most metrics. While the standard DP models offer improvement over the parametric model, they fall short of the high standards set by EDDPM. In practice, we recommend evaluating the predictive performance of these models, as illustrated in Section 5.
5 Analysis of HF-ACTION Randomized Clinical Trial: Investigating the Effect of Exercise Training in Patients with Chronic Heart Failure
We analyze the HF-ACTION trial introduced in Section 1.1 using our proposed principal stratification framework. HF-ACTION randomized medically stable outpatients with chronic systolic heart failure to usual care alone versus usual care plus a structured aerobic exercise-training program (O’Connor et al., 2009). In our analysis, the recurrent outcome is the sequence of all-cause hospitalizations over follow-up, while death is treated as an absorbing terminal event that stops further observation of subsequent hospitalizations. This setting is typical for treating patients with chronic heart failure as the cumulative burden and tempo of rehospitalizations are central measures of morbidity, yet they are intertwined with survival because longer survival increases time at risk for hospitalization while death truncates the recurrent process.
5.1 Principal stratification analysis in continuous time
Our analysis targets causal effects on rehospitalization while holding survivorship fixed by restricting attention to always-survivor principal stratum. Specifically, our primary causal estimand is the survivor-average number of recurrence ratio,
$\mathrm{SANR}(t;r)=\mu^{1}(t;r)\big/\mu^{0}(t;r),$
and $\mu^{z}(t;r)=\mathbb{E}\!\left\{N_{i}^{z}(t)\mid\mathcal{AS}(r)\right\}$. Here $z=1$ indicates exercise training and $z=0$ indicates usual care. Thus, $\mathrm{SANR}(t;r)<1$ means that among patients who would be alive up to time $r$ regardless of assignment, exercise training reduces the expected cumulative hospitalization burden up to time $t$. The estimand $\mathrm{SANR}(t;r)$ answers a counterfactual question directly relevant to the burden of heart failure management: if we could follow the same set of patients who would remain alive through time $r$ under both strategies, how would prescribing structured exercise change the expected number of hospitalizations accumulated by time $t$? Having defined the causal estimand of interest, we turn to the modeling strategies used to estimate it and assess their empirical fit. Specifically, we compare the four models considered in Section 4 using the log pseudo marginal likelihood (LPML) (Geisser and Eddy, 1979). The LPML is a Bayesian model-fit criterion derived from leave-one-out (LOO) predictive assessments of the data. The LPML values for LM, DPM, DDPM, and EDDPM are 20913.96-20913.96, 17689.37-17689.37, 15712.22-15712.22, and 15458.95-15458.95, respectively. Since a higher LPML indicates a better predictive fit in terms of LOO predictive densities, the EDDPM emerges as the superior model among the four models. We therefore present analysis results under EDDPM (5) with weakly informative conjugate priors in Section 4. Throughout this subsection, we report results for the cross-world dependence sensitivity parameter $\rho=0.5$. This is a neutral assumption in the absence of external knowledge and sensitivity analysis will be reported in Section 5.2.
We summarize the posterior distribution of $\mathrm{SANR}(t;r)$ through two informative one-dimensional slices that clarify both the time-evolution of morbidity effects and their dependence on the survivorship index $r$. Figure 1 plots $\mathrm{SANR}(t;r)$ along the traditional diagonal slice $r=t$, which corresponds to the conventional “single-indexed” always-survivor definition (e.g., Lyu et al., 2023). Figure 2 instead fixes $r\in\{720,1080,1440\}$ (i.e., 2, 3, and 4 years) and plots $\mathrm{SANR}(t;r)$ as $t$ increases within the same principal stratum, which is the key interpretive advantage of the proposed double-index formulation. Three primary findings emerge. First, posterior means suggest that adding structured exercise leads to a modest but increasing reduction in cumulative hospitalization burden over time. Across the plotted horizons, posterior mean $\mathrm{SANR}(t;r)$ is consistently below 1, indicating fewer expected hospitalizations under structured training among always-survivors. Along the diagonal $r=t$ (Figure 1), the posterior mean declines from 0.957 at $t=360$ days (1 year) to 0.90 by $t=1440$ days (4 years), corresponding to a 4.3% reduction at 1 year and a 10.1% reduction by 4 years in rehospitalization among always-survivors. The fixed-$r$ plots show a similar gradual decline: for example, when $r=1440$, the posterior mean decreases from about 0.935 at $t=360$ to 0.899 by $t=1440$ (Figure 2), suggesting a steady, cumulative reduction in the long-term burden of rehospitalization.
Second, under the traditional $r=t$ stratum definition, the cumulative benefit becomes clearly detectable only with sufficiently long follow-up. Although posterior means remain below 1 throughout, uncertainty is substantial in early-to-mid follow-up. As illustrated in Figure 1, the 95% posterior band includes the null value $\mathrm{SANR}=1$ (equivalently, $\log\mathrm{SANR}=0$) through approximately $t\approx 800$ days, and the upper bound falls below 1 only after roughly $t\gtrsim 900$ days. Thus, under the conventional time-indexed always-survivor definition, the result of HF-ACTION supports the burden-reduction benefit of structured training but only at longer horizons. This pattern is consistent with the broader understanding that the benefits of supervised exercise and other lifestyle interventions often accumulate gradually over time rather than appearing immediately among heart failure patients (Sachdev et al., 2023).
Third, fixing the survivorship stratum reveals meaningful heterogeneity that is obscured under the conventional $r=t$ perspective. The fixed-$r$ panels in Figure 2 show that the strength of evidence can differ across always-survivor strata defined by time limit $r$ even at the same time horizon $t$. In particular, at $t=800$ days the 95% posterior interval for $\mathrm{SANR}(t;r)$ still includes the null when the stratum is defined at $r=1080$ (middle panel), whereas the corresponding interval lies entirely below 1 when the stratum is defined at $r=1440$ (right panel). In other words, at the same disease burden horizon ($t=800$), restricting attention to the subgroup of patients who would survive to 1440 days under either strategy yields stronger evidence that supervised training reduces rehospitalization than expanding to the broader always-survivor stratum of those who would survive to 800 days under either treatment allocation. This distinction is substantively important for clinical practice: $\mathcal{AS}(1440)$ is a distinct, nested patient subgroup relative to $\mathcal{AS}(1080)$, and the estimated treatment effect on rehospitalization can vary across these latent risk strata. In contrast, this nuanced pattern will not be detected under the traditional $r=t$ definition, because the target subgroup changes with $t$, making it difficult to disentangle time evolution of morbidity effects from changes in the underlying principal stratum.
Overall, traditional principal stratification for censored time-to-event outcomes is often implemented with a single time index, effectively setting the survivorship index equal to the outcome horizon ($r=t$), so that the target subpopulation is $\mathcal{AS}(t)$. In recurrent-event settings, as illustrated by HF-ACTION, an apparent time trend with the single-index convention in $\mathrm{SANR}(t;t)$ can conflate (i) genuine accumulation of treatment differences in recurrent morbidity with (ii) changes in the composition of the always-survivor subgroup as $t$ increases. Our double-index formulation decouples these roles: $r$ fixes a survivorship-defined subgroup $\mathcal{AS}(r)$, while $t$ indexes the cumulative recurrent-event burden within that fixed subgroup. Consequently, plots of $\mathrm{SANR}(t;r)$ as a function of $t$ at fixed $r$ directly describe how disease burden accrues over time for the same latent subgroup, and comparisons across $r$ at a fixed $t$ reveal heterogeneity across the always-survivor stratum under different limits that are not accessible under the traditional $r=t$ definition.
Figure 1: Posterior mean and 95% posterior interval for $\mathrm{SANR}(t;r)$ with $r=t$, $\rho=0.5$.
Figure 2: Posterior mean and 95% posterior interval for $\mathrm{SANR}(t;r)$ as a function of $t$ within fixed always-survivor stratum $\mathcal{AS}(r)$ with $r\in\{720,1080,1440\}$, $\rho=0.5$.
In addition, Figure 3 reports the posterior mean and $95\%$ posterior interval for the always-survivor rate $p(r<D_{i}^{0},r<D_{i}^{1})$ in HF-ACTION, which quantifies the fraction of participants who would survive beyond $r$ under both treatment assignments. The estimated always-survivor rate decreases steadily over time, from approximately $0.94$ at $r=360$ days to approximately $0.77$ at $r=1440$ days, with a modest widening of the posterior interval at longer horizons. Two implications for heart failure care follow. First, even at multi-year horizons the always-survivor stratum remains substantial, so the principal-stratum estimands pertain to a large, clinically relevant portion of the cohort rather than to a diminishing small subgroup. Second, the downward trend reflects the nontrivial mortality in HF-ACTION, and the gradual increase in uncertainty at later time $r$ underscores that cross-world survivorship membership becomes progressively harder to identify as the horizon $r$ grows, motivating explicit sensitivity analysis in the weakly identified cross-world component.
Figure 3: Estimated always-survivor (AS) rate $p(r<D_{i}^{0},r<D_{i}^{1})$ over time (days) in HF-ACTION. Points connected by the solid line indicate the posterior mean AS rate at each time index; the shaded band denotes the corresponding $95\%$ posterior interval.
5.2 Sensitivity analysis for cross-world dependence
This section complements the empirical results above by assessing robustness to the cross-world frailty correlation parameter $\rho$, which governs dependence between the two potential processes and is not identified from the observed data. Figures 4 summarize posterior inference for $\mathrm{SANR}(t;r)$ under five values $\rho\in\{0.1,0.3,0.5,0.7,0.9\}$. We focus on a positive grid because the scientifically plausible concern in this setting is positive cross-world dependence induced by shared latent health status, whereas negative dependence appears much less credible. For each $\rho$, we display both the diagonal slice $r=t$ and fixed-stratum slices with $r\in\{720,1080,1440\}$ in Figure 4. For $\rho\in\{0.1,0.3,0.5,0.7\}$, the qualitative pattern is stable: posterior means are below 1 and generally decline with $t$, indicating a consistent direction of effect toward reduced cumulative hospitalization burden under exercise training among always-survivors. Varying $\rho$ between $0.1$ and $0.7$ does not shift the level of the curves and alters posterior uncertainty, thus the substantive interpretation of the HF-ACTION results remains the same. For example, under $\rho=0.3$, posterior means of $\mathrm{SANR}(t;r)$ remain below 1 across the plotted horizons and generally decline with $t$, indicating a persistent, modest reduction in cumulative hospitalization burden under exercise training among always-survivors. Along the diagonal slice, the posterior mean decreases from $0.960$ at $t=360$ days to $0.902$ by $t=1440$ days. The $95\%$ posterior interval continues to include the null value $\mathrm{SANR}=1$ throughout $t=1080$, which is a slightly longer horizon than the primary analysis under $\rho=0.5$, but the overall trend remains similar. The fixed-$r$ summaries also show a similar pattern to the primary analysis.
(a) $\rho=0.1$
(b) $\rho=0.3$
(c) $\rho=0.5$
(d) $\rho=0.7$
(e) $\rho=0.9$
Figure 4: Posterior mean and $95\%$ posterior interval for $\mathrm{SANR}(t;r)$ in HF-ACTION under five sensitivity parameters $\rho$. In each row, the left panel shows the diagonal slice $r=t$, and the right panel shows fixed always-survivor stratum $r\in\{720,1080,1440\}$.
However, a somewhat different pattern emerges under the more extreme choice of $\rho=0.9$. A large value of $\rho$ implies very strong cross-world correlation between the frailties $\gamma^{0}$ and $\gamma^{1}$, so that the model effectively approximates a single-frailty specification with $\gamma^{0}\approx\gamma^{1}$, similar in spirit to the single shared-frailty formulation adopted in Lyu et al. (2023) and Comment et al. (2025). Under this more extreme assumption, the $95\%$ posterior interval continues to include the null value $\mathrm{SANR}=1$ throughout the displayed range, with the upper bound remaining close to 1 even at the longest follow-up times. Thus, relative to the primary dual-frailty analysis, imposing near-perfect cross-world dependence yields more conservative inference; although the posterior means remain consistently below 1, the evidence for a reduction in rehospitalization burden becomes less decisive. The fixed-$r$ summaries show the same pattern. For each survivorship stratum $r$, the posterior means remain below 1 and decline gradually with $t$, but the $95\%$ posterior intervals typically continue to overlap 1 over the observed follow-up period. In particular, at $t=900$ days, the posterior interval still includes $\mathrm{SANR}=1$ both when the always-survivor stratum is defined at $r=1080$ and when it is defined at $r=1440$. Hence, under $\rho=0.9$, the stratum-specific separation from the null that was visible under the dual-frailty analysis with $\rho=0.1$ to $0.7$ is largely attenuated.
Taken together, these results indicate that the qualitative direction of the estimated exercise effect among always-survivors is robust, in the sense that the posterior mean $\mathrm{SANR}$ remains below 1 across the sensitivity analysis. When $\rho$ lies in the moderate range $\rho=0.1$ to $0.7$, the dual-frailty formulation reveals meaningful stratum-specific evidence of reduced rehospitalization burden. By contrast, when $\rho=0.9$, the near single-frailty assumption suppresses that signal and pulls the inference toward the null. This attenuation is substantively important; in our setting, treatment may affect rehospitalization risk and mortality risk through partially distinct latent pathways, so forcing the two cross-world frailties to be nearly identical runs the risk of masking clinically meaningful heterogeneity. For this reason, the dual-frailty framework is not only a flexible generalization of prior shared-frailty approaches, but also an important modeling device for recovering treatment-effect patterns that would otherwise be obscured under an overly restrictive dependence assumption.
@@ 6 Concluding Remarks
Motivated by HF-ACTION, a real-world randomized clinical trial where the endpoint of the number of rehospitalizations is clinically important but subject to truncation by death, we proposed a new principal stratification approach for causal recurrent event analysis. Our work makes two main contributions. First, we introduce a new principal causal estimand for recurrent events,together with a partial identification result that enables transparent sensitivity analysis for the cross-world dependence structure of potential outcomes. Second, we develop the EDDPM prior, a nested Bayesian nonparametric construction that jointly models recurrent and terminal events while embedding covariate dependence and naturally accommodating the frailty structure required for identification. By separating terminal-event clusters from finer recurrent-event subclusters, the EDDPM addresses the unfavorable random partition behavior of standard DP priors, and offers flexibility over parametric approaches (Nevo and Gorfine, 2022; Lyu et al., 2023). We provided a fully tractable Gibbs sampling algorithm for posterior inference and demonstrated through simulation studies that the proposed method outperforms traditional DP-based and parametric approaches in both accuracy and robustness.
Beyond these methodological contributions, the HF-ACTION analysis also illustrates why the proposed framework is useful compared to existing alternatives in practice. Above all, the double-index estimand $\mathrm{SANR}(t;r)$ yields different and richer conclusions than the conventional single-index alternative. Because enforcing $r=t$ makes the always-survivor subgroup shrink as the outcome horizon advances, any apparent time trend inevitably mixes genuine changes in rehospitalization dynamics with compositional shifts in the at-risk population, whereas decoupling $r$ from $t$ holds the target subpopulation fixed and reveals useful heterogeneity across always-survivor strata that is entirely invisible under the traditional definition. Second, our sensitivity analysis approach for the cross-world frailty correlation $\rho$ shows that the choice between a dual-frailty and a single-frailty specification is not merely a matter of modeling generality but can change the substantive conclusions drawn from the data. That is, at moderate values of $\rho$ the dual-frailty formulation detects a gradually accumulating reduction in rehospitalization burden that eventually separates from the null, whereas forcing the two frailties to be nearly identical (as assumed in existing shared-frailty approaches in Lyu et al. (2023)) attenuates this signal. Taken together, these two features of the analysis highlight the applied value of the proposed framework for chronic-disease trials in which recurrent morbidity and mortality are intertwined. Our HF-ACTION case study encourages practitioners to move beyond single-index, single shared-frailty analyses when the scientific question concerns cumulative disease burden among always-survivors.